OpenStudy (hysenberg):

how to evaluate roots of (1)^(1/8)

OpenStudy (zzr0ck3r):

\(1^x=1\) for any number \(x\)

zepdrix (zepdrix):

8th roots of unity? You can use them complex number shinanigans and De'Moivre's Theorem to help you out.

OpenStudy (zzr0ck3r):

Sorry, assumed real...

zepdrix (zepdrix):

\[\large\rm 1=e^{0+2k \pi i}\]\[\large\rm 1^{1/8}=e^{\frac{2k \pi i}{8}}=e^{\frac{k \pi i}{4}}\]Consider the fist 8 integer values of k, starting from 0: k=0,1,2,3,4,5,6,7. These will give us all our unique 8th roots of 1. \(\large\rm e^{0}=1\) \(\large\rm e^{\pi i/4}\) I guess you'll have to convert to sines and cosines to simplify some of these roots :)

zepdrix (zepdrix):

I can go into more detail if you need, looks like you went offline for now though :D